# Geometry-Aware Edge Pooling for Graph Neural Networks

### v1

@@ B.3.3 Edge contraction on graphs
We will first recall important results about magnitude in the context of (strictly) positive definite finite metric spaces and refer the interested reader to Leinster [28] for further details.
Proposition 1 (Proposition 2.4.3, [28]).
@@
$\text{Mag}(X)=\sup_{v\in\mathbb{R}^{n}\backslash\{0\}}\frac{(\sum_{i=0}^{n}v_{i})^{2}}{v^{t}\zeta_{X}v}.$ (8)
Proposition 1 implies that the magnitude of positive definite metric spaces is always positive. Another consequence of this result is a monotonicity property on subsets of these metric spaces.
Corollary 2 (Corollary 2.4.4, [28]).
@@
Finally, by Proposition 1 and Inequality 10, we get that
$\begin{split}\text{Mag}(G_{1})&=\sup_{v\in\mathbb{R}^{n}\backslash\{0\}}\frac{(\sum_{i=0}^{n}v_{i})^{2}}{v^{t}\zeta_{X_{1}}v}\geq\sup_{v\in\mathbb{R}^{m}\backslash\{0\}}\frac{(\sum_{i=0}^{n}v_{i})^{2}}{v^{t}\zeta_{X_{1}}v}\\ &\geq\sup_{v\in\mathbb{R}^{m}\backslash\{0\}}\frac{(\sum_{i=0}^{m}v_{i})^{2}}{v^{t}\zeta_{X_{2}}v}=\text{Mag}(G_{2})\end{split}$
@@ B.3.4 Bounding magnitude by spread
Theorem 6 (Theorem 2.2. [41]).
@@
We will now use this bound as well as the results in Section B.3.3 to investigate the relationship between MagEdgePool and SpreadEdgePool. Through a process of iterated edge contraction, our pooling algorithm produces a sequence of hierarchically pooled graphs (as described in Section 3.1 and Section C.4). Note that it is not guaranteed that MagEdgePool and SpreadEdgePool yield the same sequence. For this reason, we will refer to the graphs resulting from the $k^{th}$ edge-contraction with magnitude and spread by $G^{(k)}$ and $\widetilde{G}^{(k)}$ respectively.
For each $k$ the edge contraction map $G^{(k)}\rightarrow G^{(k+1)}$ is a surjection on the underlying vertex sets $X^{(k)}$ and $X^{(k+1)}$ respectively, i.e $X^{(k+1)}\subset X^{(k)}$. Moreover, we will assume that for any $k$ this map is distance-decreasing. That is, $d^{(k+1)}(f(x_{i}),f(x_{j}))\leq d^{(k)}(x_{i},x_{j})$ for all $x_{i},x_{j}\in X^{(k+1)}$.
Recall that for any $k$, $\text{Mag}(G^{(k)})$ is the magnitude of a finite positive definite metric space $(X^{(k)},d^{(k)})$. Then, by Theorem 4, we deduce that for any $k$,
@@
Note that by construction, scoring the edges in Algorithm C.4 translates into the following:
@@
Then, by monotonicity of magnitude (Inequality 11),
@@
Assume that $|\text{Mag}(G^{(k-1)})-\text{Sp}(G^{(k)})|\leq C\Delta^{(k)}\text{Sp}(G)$ for some constant $C>0$. By the monotonicity of magnitude (Theorem 4), we get that $\text{Mag}(G^{(k)})\leq\text{Mag}(G^{(k-1)})$ and,
@@ C.1 Hardware and Software
The experiments reported in our study were implemented using spektral 1.3.1 [24]***https://graphneural.network/ available under an MIT license., and tensorflow 2.16.2 [1]https://pypi.org/project/tensorflow/2.16.2/ available under the Apache Software License (Apache 2.0).. As further detailed in Section C.5.2 and Table 1, we base our graph classification experiments on the benchmark setup and code by Grattarola et al. [25]https://github.com/danielegrattarola/SRC available to the research community (Grattarola et al. [25])., which also include implementations for the pooling layers compared across our study. By relying on this existing framework, we aim to ensure the reproducibility of our results.
Further, to calculate magnitude and spread we rely on magnipy, a Python package by Limbeck et al. [31]§§§https://github.com/aidos-lab/magnipy available under a BSD 3-Clause License. for magnitude and diversity computations. Further, as a novel contribution of this paper, we extend the computation of magnitude to graph data and novel graph metrics. Specifically, we modify the computations, so that the magnitude of disconnected subgraphs is computed separately (based on Theorem 2) using the NetworkXhttps://github.com/networkx/networkx available under a BSD 3-Clause License. package. We also implement graph distances that have not previously been used to compute magnitude, such as the diffusion distances detailed in Section 2.3. Further details on the code for implementing our proposed pooling methods and our experiments can be found in our supplementary code submission.
All experiments were conducted on a high performance cluster with hardware specifications as detailed in Table 2. In particular, all experiment were run requesting one GPU with 32GB video memory or less.
Table 2: Summary of the compute resources used for our experiments.
@@ C.2 Datasets
We briefly describe the graph datasets analysed throughout our work. Simulated graphs, as used for Figure 1, are created using either PyGSP https://pygsp.readthedocs.io/en/stable/ available under a BSD-3-Clause license. or NetworkX******https://github.com/networkx/networkx available under a BSD 3-Clause License. and all example graphs were created to consist of 64 nodes.
For our main graph classification experiments, we analyse six graph datasets taken from biological or chemical applications [40, 10, 37, 17, 38], and two datasets which represent social networks [43]. All datasets are taken either from the TUDataset††††††https://chrsmrrs.github.io/datasets/ available under a CC BY 4.0 license. benchmark [35] or the Open Graph Benchmark‡‡‡‡‡‡https://ogb.stanford.edu/ available under an MIT licence..
More specifically, the main results in Table 1 analyse the following graph classification datasets described in Table 3. Note that we only consider node and not edge features for our experiments.
Table 3: Summary of the graph datasets considered for our experiments.
@@ C.3 Magnitude and Spread Computations
Across our experiments, we compute magnitude and spread as outlined in the main text, implemented in our code submission, and further described in Section C.1. Elaborating on these descriptions, we now aim to give an extended explanation of practical and theoretical considerations for computing the magnitude of graphs in practice.
@@
In mathematical literature, the magnitude of graphs is often studied with the shortest path metric [29]. However, shortest path distances are not guaranteed to be of negative type, thus leading to scenarios and well-known examples for which the similarity matrix is not invertible and magnitude based on this metric cannot be computed [28]. In comparison, resistance distances, diffusion distances, or Euclidean distance always permit the computation of magnitude. Because of this difference in the choice of distance metric, we note that our definition of the magnitude of a graph in Section 2.2 differs from the definition used by Leinster [29].
@@
Note that magnitude and spread can also be defined as multi-scale functions i.e. $t\mapsto\text{Mag}((X,t\cdot d))$ for a metric space $(X,d)$ and a scale parameter $t\in\mathbb{R}^{+}$. This parameter $t$ can be likened to choosing a kernel bandwidth or the scale of distances or similarity determining when observations are considered to be distinct. In practical applications, it is advisable to carefully consider the choice of scaling factor $t$ or the type of normalisation used to compare distances [31]. Limbeck et al. [31] propose a heuristic that uses root-finding to find a suitably large $t$. However, this requires repeated computations of magnitude and increases the computational costs. A faster and more desirable default choice of $t$ would be based solely on the distance metric. For distances that are not otherwise scaled or normalised, we therefore recommend the usage of faster heuristics, such as the median heuristic for choosing the kernel-bandwidth i.e. the scale parameter $t$ [22]. For diffusion distances, we find that setting $t=1$ is sufficient for our goals. This is because, as discussed in Section 2.3, diffusion distances are computed from the normalised graph Laplacians and are inherently comparable across graphs. Nevertheless, investigating magnitude and spread as multi-scale functions on graphs remains an interesting extension for further work.
@@ C.4 Pooling Algorithm
We now detail our pooling algorithm introduced in Section 3.1 by describing a pseudocode implementation. Note that to describe the algorithm we assume we have pre-selected a distance metric for computing either magnitude or spread.
@@
0:  input graph $G=(X,E)$, node features $\mathbf{F}\in\mathbb{R}^{n\times d}$, pooling ratio $r\in(0,1]$, diversity measure $\text{Mag}(G)\text{ or }\text{Sp}(G)$
@@
2:  Initialise the set of contracted edges $E_{c}\leftarrow\emptyset$
@@
18:  for each supernode representative $w\in\mathcal{S}$ do
@@
20:     Compute aggregated feature:
@@ C.5 Extended Experimental Details
C.5.1 Structure Preservation Experiment and Pooling Ratios
For reporting the structure preservation results described in Section 4.2, we considered multiple different proposals for what it means to preserve graph structure during pooling. In the end, we settled to compare the spectral distance between the symmetrically normalised graph Laplacians as an established measure of spectral property preservation, and investigated the relative difference in magnitude between the original graph $G$ and the pooled graph $G^{\prime}$ computed from diffusion distances. Specifically, the reported relative magnitude difference is calculated as
@@ C.5.1 Structure Preservation Experiment and Pooling Ratios
For this experiment we further varied the pooling ratios across different pooling methods. However, some of the pooling layers considered in our study, specifically NDP, Graclus and NMF, were configured to always pool graphs to around half their size. To allow us to compare these methods across increasing pooling ratios, we choose to reapply these pooling operations repeatedly, which is why these three pooling methods are only evaluated at pooling ratios that are powers of 0.5.
C.5.2 Graph Classification Experiment
Our graph classification architecture follows the experimental setup described in Table 1 and is based the benchmark by Grattarola et al. [25]. Across our main classification experiment detailed in Table 1, different pooling layers are configured to reduce each input graph to around 50% of the number of nodes in the original graphs. Depending on the pooling method, this is chosen so each graph is reduced to 50% of its original size $k=\lfloor 0.5*N\rfloor$, (for NDP, Graclus, MagEdgePool, SpreadEdgePool, TopKPool, and SAGPool), or to 50% the average size of all graphs in the training dataset $k=\lfloor 0.5*\bar{N}\rfloor$ (for DiffPool, and MinCutPool). Interpreting the experimental results in Table 1 it is thus of interest that the sizes of the pooled graphs can vary across pooling layers, which might explain some of the difference in performance between the fixed-size methods DiffPool and MinCUT compared to more adaptive pooling methods.

### v2

@@ B.3.3 Edge contraction on graphs
We first recall important results about magnitude in the context of (strictly) positive definite finite metric spaces and refer the interested reader to Leinster [37] for further details.
Proposition 1 (Proposition 2.4.3, Leinster [37]).
@@
[…]
˜1 implies that the magnitude of positive definite metric spaces is always positive. Another consequence of this result is a monotonicity property on subsets of these metric spaces.
Corollary 2 (Corollary 2.4.4, Leinster [37]).
@@
Finally, by ˜1 and Equation˜10, we get that
[…]
@@ B.3.4 Bounding magnitude by spread
Theorem 6 (Theorem 2.2. from Willerton [52]).
@@
We will now use this bound as well as the results in Section˜B.3.3 to investigate the relationship between MagEdgePool and SpreadEdgePool. Through a process of iterated edge contraction, our pooling algorithm produces a sequence of hierarchically pooled graphs (as described in Section˜3.1 and Section˜C.4). Note that it is not guaranteed that MagEdgePool and SpreadEdgePool yield the same sequence. For this reason, we will refer to the graphs resulting from the $k^{th}$ edge-contraction with MagEdgePool and SpreadEdgePool by $G^{(k)}$ and $\widetilde{G}^{(k)}$ respectively.
For each $k$, the edge contraction map $G^{(k)}\rightarrow G^{(k+1)}$ is a surjection on the underlying vertex sets $X^{(k)}$ and $X^{(k+1)}$ respectively, i.e $X^{(k+1)}\subset X^{(k)}$. Moreover, we will assume that for any $k$ this map is distance-decreasing. That is, $d^{(k+1)}(f(x_{i}),f(x_{j}))\leq d^{(k)}(x_{i},x_{j})$ for all $x_{i},x_{j}\in X^{(k+1)}$.
Recall that for any $k$, $\text{Mag}(G^{(k)})$ is the magnitude of a finite positive definite metric space $(X^{(k)},d^{(k)})$. Then, by ˜4, we deduce that for any $k$,
@@
Note that by construction, scoring the edges in Algorithm˜1 translates into the following:
@@
Then, by the monotonicity of magnitude, i.e. Equation˜11,
@@
Assume that $|\text{Mag}(G^{(k-1)})-\text{Sp}(G^{(k)})|\leq C\Delta^{(k)}\text{Sp}(G)$ for some constant $C>0$. By the monotonicity of magnitude (˜4), we get that $\text{Mag}(G^{(k)})\leq\text{Mag}(G^{(k-1)})$ and,
@@
B.4 Expressivity
While studying the expressive power of GNNs, we aim to evaluate their ability to generate different outputs for non-isomorphic graphs. Here, we will analyse the expressive power of our pooling methods MagEdgePool and SpreadEdgePool within the Select-Reduce-Connect framework introduced by Grattarola et al. [32] for describing pooling operators. Let $G=(X,E)$ be a graph and $\mathbf{F}\in\mathbb{R}^{|X|\times f}$ be the node features associated to the nodes in $G$. Then, a graph pooling operator is regarded as a function $\text{{POOL}}\colon(\mathbf{F},G)\mapsto(\mathbf{F}_{P},G_{P})$ where $\mathbf{F}_{P}$ denotes the pooled node features and $G_{P}=(X_{P},E_{P})$ the pooled graph with $|X_{P}|\leq|X|$. Pooling is described as a combination of three elementary functions: selection (SEL), reduction (RED), and connection (CON). The selection function clusters the nodes of the input graph into super-nodes, so that $\text{{SEL}}\colon G\mapsto\mathcal{S}=\{\mathcal{S}_{j}\}_{j=1}^{|X_{P}|}$ where $\mathcal{S}_{j}=\{S_{ij}\}_{i=1}^{|X|}$ and $S_{ij}$ is the membership score of the node $i$ to super-node $j$. Node selection can be represented as the assignment matrix $S\in\mathbb{R}^{|X|\times|X_{P}|}$ with entries $S_{ij}$. Based on this selection, the reduction function aggregates node features of all nodes that are assigned to the same super-node, i.e. $\textbf{RED}\colon(\mathbf{F},S)\mapsto\mathbf{F}_{P}$. Then, the connection function, CON, generates the edges and determines the connectivities between super-nodes in the pooled graph. Finally, to study expressivity, we note that hierarchical graph pooling is typically applied in GNN architectures after some initial message passing layers. Let $G^{L}$ denote the graph resulting from a block of $L$ MP layers and $\mathbf{F}^{L}\in\mathbb{R}^{|X|\times f}$ the corresponding feature matrix [8].
Theorem 7.
(Theorem 1 from Bianchi and Lachi [8]) Let $G_{1}=(X_{1},E_{1})$ and $G_{2}=(X_{2},E_{2})$ be two graphs equipped with node features $\mathbf{F}_{i}\in\mathbb{R}^{|X_{i}|\times f}$ for $i=1,2$. Assume that $G_{1}\neq_{\text{WL}}G_{2}$, i.e. that $G_{1}$ and $G_{2}$ are distinguishable by the Weisfeiler-Leman isomorphism test. Apply a block of L MP layers to get $G^{L}_{1}$ and $G^{L}_{2}$ as well as $\mathbf{F}^{L}_{1}$ and $\mathbf{F}^{L}_{2}$. Let POOL be a pooling operator placed after these MP layers to get $G_{1_{P}}=\text{{POOL}}(G^{L}_{1})$ and $G_{2_{P}}=\text{{POOL}}(G^{L}_{2})$ associated with the node features $\mathbf{F}_{1_{P}}$ and $\mathbf{F}_{2_{P}}$ in $\mathbb{R}^{k\times f}$. Then, $G_{1_{P}}$ and $G_{2_{P}}$ will have different node features (up to permutation) provided the following conditions hold:
  1. 1.
    $\sum_{i=1}^{|X_{1}|}\mathbf{F}^{L}_{{1}_{[i,:]}}\neq\sum_{i=1}^{|X_{2}|}\mathbf{F}^{L}_{{2}_{[i,:]}}$,
  2. 2.
    The memberships generated by SEL satisfy $\sum_{j=1}^{k}S_{ij}=\lambda$, with $\lambda>0$ for each node $i$, i.e., the cluster assignment matrix S is a right stochastic matrix up to the global constant $\lambda$,
  3. 3.
    The reduction function satisfies $\text{{RED}}\colon(\mathbf{F}^{L},S)\mapsto\mathbf{F}_{P}=S^{T}\mathbf{F}^{L}$.
The WL test will identify that two graphs with different multisets of node features are non-isomorphic based on the injectivity of the colouring function of the WL algorithm. ˜7 then guarantees that $G_{1_{P}}\neq_{WL}G_{2_{P}}$ thus ensuring that the pooling operation POOL preserves expressivity.
Corollary 3.
MagEdgePool and SpreadEdgePool satisfy the sufficient conditions outlined in ˜7 when using sum aggregation for pooling the node features.
Proof.
Condition 1 is independent to the choice of pooling layer and instead relates to the expressivity of MP layers. It is guaranteed for any MP layer that is as powerful as the 1-WL test [8, 25]. Conditions 2 and 3 hold trivially by construction. Each super-node is the result of edge contractions and node features are aggregated via summation. Hence, each vertex is assigned to a unique super-node and the selection matrix is constructed as $S_{ij}=1$ if the node $i$ is contained in super-node $j$ and $S_{ij}=0$ otherwise. This ensures that $\sum_{j=1}^{k}S_{ij}=1$ for every node $i$ fulfilling condition 2. The resulting SEL function can be represented using the cluster assignment matrix $S$ obtained as a product of these elementary (contraction) operations represented by matrices. This yields the RED function described in our algorithm in Algorithm˜1 as a map $\mathbf{F}^{L}\mapsto S^{T}\mathbf{F}^{L}$ that respects condition 3. ∎
@@ C.1 Hardware and Software
The experiments reported in our study were implemented using spektral 1.3.1 [31] [^2] https://graphneural.network/ available under an MIT license., and tensorflow 2.16.2 [1] [^3] https://pypi.org/project/tensorflow/2.16.2/ available under the Apache Software License (Apache 2.0).. As further detailed in Section˜C.5.3 and Section˜4.1, we base our graph classification experiments on the benchmark setup and code by Grattarola et al. [32] [^4] https://github.com/danielegrattarola/SRC available to the research community (Grattarola et al. [32])., which also include implementations for the pooling layers compared across our study. By relying on this existing framework, we aim to ensure the reproducibility of our results.
Further, to calculate magnitude and spread we rely on magnipy, a Python package by Limbeck et al. [40] for magnitude and diversity computations. [^5] https://github.com/aidos-lab/magnipy available under a BSD 3-Clause License. Further, as a novel contribution of this paper, we extend the computation of magnitude to graph data and novel graph metrics. Specifically, we modify the computations, so that the magnitude of disconnected subgraphs is computed separately (based on ˜2) using the NetworkX [^6] https://github.com/networkx/networkx available under a BSD 3-Clause License. package. We also implement graph distances that have not previously been used to compute magnitude, such as the diffusion distances detailed in Section˜2.3. Further details on the code for implementing our proposed pooling methods and our experiments can be found in our supplementary code submission as well as on GitHub. Finally, we publish a reproducible PyTorch implementation of our pooling methods as mag_edge_pool [^7] https://github.com/aidos-lab/mag_edge_pool available under a BSD 3-Clause License. on GitHub.
All experiments were conducted on a high-performance cluster with hardware specifications as detailed in Table˜S.1. In particular, all experiment were run requesting a single GPU with 32 GB video memory or less.
Table S.1: Summary of the compute resources used for our experiments.
@@ C.2 Datasets
We briefly describe the graph datasets analysed throughout our work. Simulated graphs, as used for Figure˜1, are created using either PyGSP  [^8] https://pygsp.readthedocs.io/en/stable/ available under a BSD-3-Clause license. or NetworkX [^9] https://github.com/networkx/networkx available under a BSD 3-Clause License. and all example graphs were created to consist of 64 nodes.
For our main graph classification experiments, we analyse six graph datasets taken from biological or chemical applications [49, 11, 46, 22, 47], and two datasets which represent social networks [54]. All datasets are taken either from the TUDataset [^10] https://chrsmrrs.github.io/datasets/ available under a CC-BY-4.0 license. benchmark [45] or the Open Graph Benchmark [^11] https://ogb.stanford.edu/ available under an MIT licence..
More specifically, the results in Table˜1, Table˜2, and Table˜S.8 analyse the following graph datasets described in Table˜S.2. Note that we only consider node and not edge features for our experiments.
Table S.2: Summary of the graph datasets considered for our experiments.
@@
OGBG-MOlESOL OGB regression 9 1128 13 14
OGBG-MOlFREESOLV OGB regression 9 642 9 8
OGBG-MOLLIPO OGB regression 9 4200 27 30
@@ C.3 Magnitude and Spread Computations
Across our experiments, we compute magnitude and spread as outlined in the main text, implemented in our code submission, and further described in Section˜C.1. Elaborating on these descriptions, we now aim to give an extended explanation of practical and theoretical considerations for computing the magnitude of graphs in practice.
@@
In mathematical literature, the magnitude of graphs is often studied with the shortest path metric [38]. However, shortest path distances are not guaranteed to be of negative type, thus leading to scenarios and well-known examples for which the similarity matrix is not invertible and magnitude based on this metric cannot be computed [37]. In comparison, resistance distances, diffusion distances, or Euclidean distance always permit the computation of magnitude. Because of this difference in the choice of distance metric, we note that our definition of the magnitude of a graph in Section˜2.2 differs from the definition used by Leinster [38]. While we choose to investigate diffusion distances, we note that the distance metric can easily be replaced if needed to explore alternative geometries.
Diffusion geometry.
For further research, we believe that the usage of diffusion distances offers the chance to leverage a rich theory on approximation methods via landmarks [42], or localised diffusion computations [16], which can lead to further computational improvements and extension of our methods.
@@
Note that magnitude and spread can also be defined as multi-scale functions i.e. $t\mapsto\text{Mag}((X,t\cdot d))$ for a metric space $(X,d)$ and a scale parameter $t\in\mathbb{R}^{+}$. This parameter $t$ can be likened to choosing a kernel bandwidth or the scale of distances or similarity determining when observations are considered to be distinct. In practical applications, it is advisable to carefully consider the choice of scaling factor $t$ or the type of normalisation used to compare distances [40]. Limbeck et al. [40] propose a heuristic that uses root-finding to find a suitably large $t$. However, this requires repeated computations of magnitude and increases the computational costs. A faster and more desirable default choice of $t$ would be based solely on the distance metric. For distances that are not otherwise scaled or normalised, we therefore recommend the usage of faster heuristics, such as the median heuristic for choosing the kernel-bandwidth i.e. the scale parameter $t$ [29]. For diffusion distances, we find that setting $t=1$ is sufficient for our goals. This is because, as discussed in Section˜2.3, diffusion distances are computed from the normalised graph Laplacians and are inherently comparable across graphs. Nevertheless, investigating magnitude and spread as multi-scale functions on graphs remains an interesting extension for further work.
@@ C.4 Pooling Algorithm
We now detail our pooling algorithm introduced in Section˜3.1 by describing a pseudocode implementation. Note that to describe the algorithm we assume we have pre-selected a distance metric for computing either magnitude or spread.
@@
0: input graph $G=(X,E)$, node features $\mathbf{F}\in\mathbb{R}^{|X|\times f}$, pooling ratio $r\in(0,1]$, diversity measure $\text{Mag}(G)\text{ or }\text{Sp}(G)$
@@
2: Initialise the set of edges adjacent to a contracted edge $E_{c}\leftarrow\emptyset$
@@
18:for each super-node representative $w\in\mathcal{S}$ do
@@
20:  Compute the aggregated features:
@@ C.5 Extended Experimental Details
C.5.1 Overview Experiment
To create the illustration in Figure˜1 and visually compare the outputs of different pooling methods, we follow the experimental setup by Grattarola et al. [32] on understanding structure preservation in graph pooling layers. Specifically, we simulate a ring graph with 64 nodes, a barbell graph with 20 nodes on each side connected by 24 nodes in the middle, and a sensor graph with 64 nodes. Graphs are then pooled to a pooling ratio of approximately $50\%$. Because some pooling methods (e.g. Graculus or NDP) do not give exact control over the number of vertices, but pool graphs to approximately half of their original size, the number of nodes visualised in Figure˜1 can vary across methods. All trainable pooling layers were then trained in a self-supervised manner to optimise the following spectral loss between the original graph $G$ and the pooled graph $G^{\prime}$:
$\mathcal{L}(G,G^{\prime})=\sum_{i=0}^{f}F^{\top}_{:,i}LF_{:,i}-F^{\prime\top}_{:,i}L^{\prime}F^{\prime}_{:,i}$ (14)
C.5.2 Structure Preservation Experiment and Pooling Ratios
For reporting the structure preservation results described in Section˜4.2, we considered multiple different proposals for what it means to preserve graph structure during pooling. In the end, we settled to compare the spectral distance between the symmetrically normalised graph Laplacians as an established measure of spectral property preservation, and investigated the relative difference in magnitude between the original graph $G$ and the pooled graph $G^{\prime}$ computed from diffusion distances. Specifically, the reported relative magnitude difference is calculated as
@@ C.5.2 Structure Preservation Experiment and Pooling Ratios
For this experiment we further vary the pooling ratios across different pooling methods. However, some of the pooling layers considered in our study (NDP, Graclus and NMF) were configured to always pool graphs to around half their size. To allow us to compare these methods across increasing pooling ratios, we choose to reapply these pooling operations repeatedly, which is why these three pooling methods are evaluated at pooling ratios that are powers of 0.5.
C.5.3 Graph Classification Experiment
Our graph classification architecture follows the experimental setup described in Section˜4.1 and is based the benchmark by Grattarola et al. [32]. Across our main classification experiment detailed in Section˜4.1, different pooling layers are configured to reduce each input graph to around 50% of the number of nodes in the original graphs. Depending on the pooling method, this is chosen so each graph is reduced to 50% of its original size $k=\lfloor 0.5*N\rceil$, (for NDP, Graclus, MagEdgePool, SpreadEdgePool, TopKPool, and SAGPool), or to 50% the average size of all graphs in the training dataset $k=\lfloor 0.5*\bar{N}\rfloor$ (for DiffPool, and MinCutPool). Interpreting the experimental results in Table˜1 it is thus of interest that the sizes of the pooled graphs can vary across pooling layers, which might explain some of the difference in performance between the fixed-size methods DiffPool and MinCUT compared to more adaptive pooling methods.
C.5.4 Graph Regression
As a further ablation study, we aim to assess whether the results reported in Table˜1 for graph classification tasks remain consistent for further graph regression tasks. To this end, we use three molecular datasets from the OGB benchmark with their predefined test, training, and validation splits [33]. Further, we adjust the GNN architecture described in Section˜4.1 and Section˜C.5.3 to use the MSE as a training loss, the RMSE for performance evaluation, a linear final activation for the readout MLP, an early stopping patience of 100 epochs, and blocks of two convolutional layers instead of single layers. Similar to before, we only use node features as inputs. Note that stronger, domain-specific, and purpose-built models exist for molecular regression tasks that utilise both atom and bond information [33]. Hence, our goal is not to reach state of the art performance. Rather, we aim to compare the performance of different pooling layers and evaluate the information loss due to the pooling operations themselves. Finally, Table˜2 reports the RMSE on the test dataset across ten repeats using varying random seeds.